# Vulnerabilities of quantum key distribution systems in visible range

### v1

@@ III.1 Optical attenuators
In the zero-voltage case, it turned out that the measured insertion losses are higher than the values reported in the datasheet. For EM VOA and EO VOA, the mimimal losses 10.7 dB and 21.8 dB are reached at the wavelengths 761 nm and 557 nm, respectively.
Referring to Figs. 2(a) and (b), when the applied voltage is 5 V (blue lines), the losses are shown to vary in the range between 13.7 dB and 30 dB for EM VOA, whereas, for EO VOA, they are ranged from 23.7 dB to 50 dB. Note that these maximum values of losses are approximate as they correspond to the noise level of our experimental setup and depend on the number of neutral filters.
Insertion losses measured in relation to the wavelength for six FOAs are shown in Figs. 2(c) and (d). For three absorption based FOAs (see Figs. 2(c)), according to the values reported by the manufacurers, the insertion losses of one FOA is 10 dB, whereas the losses for the other two FOAs are the same and equal to 20 dB. For the scattering based FOAs (see Figs. 2(d)), there are three different reported values of losses at the operating wavelength ($\lambda=1550$ nm): 5 dB, 15 dB and 20 dB. Our experimental results are in good agreement with these values.
The insertion losses of the absorption based FOAs measured in the 400-800 nm range are shown in Fig. 2(c). It can be seen that, for each FOA, there are two transmittance windows where the insertion losses are noticeably lower than the value at the operating wavelength indicated in the legend. The window ranged between 400 nm and 430 nm is located in the vicinity of local minima for the two 20 dB attenuators with $\alpha_{\mathrm{dB}}\approx 8.1$ and $\alpha_{\mathrm{dB}}\approx 8.6$ at the wavelength $405$ nm (for the 10 dB attenuator, the corresponding value is $\alpha_{\mathrm{dB}}\approx 2.7$) Similarly, the other window in the 750-800 nm wavelengths range is close to the local minima at $\lambda\approx 780$ nm with the inserion losses estimated at $\alpha_{\mathrm{dB}}\approx 3.4$ and $\alpha_{\mathrm{dB}}\approx 4.3$ and $\alpha_{\mathrm{dB}}\approx 0.9$.
Figure 2(d) shows the wavelength dependencies of the insertion losses measured for the scattering based FOAs. It is seen that, for the 20 dB attenuator, the insertion losses insignificantly change with wavelength varying from 27 dB to 25 dB. Similar conclusion applies to the 15 dB (5 dB) attenuator, where variations are between 15.6 dB (11.2 dB) and 13 dB (10.8 dB).
III.2 Wavelength-division multiplexing components
In QKD systems, wavelength-division multiplexing (WDM) components can be used as countermeasures by suppressing radiation with wavelengths outside the range utilized by the communication channel between Alice and Bob. In the simplest case, the key elements of the internal design of fiber WDM component are thin films deposited on a substrate. Ideally, at the working wavelength, light will enter the common port (the com-port) and, after passing through the thin films and the substrate, will propagate to the corresponding port of the operating wavelengths grid. Other wavelengths are directed to the reflecting port (the ref-port). This port might be connected with a watchdog detector that monitor Eve’s probing pulses. The latter assumes that the sensitivity range of the watchdog detector is broad enough to cover wavelengths of potential attacks.
In practice, there is a number of factors limiting the efficiency of the WDM-components. These include the substrate transmittance as well as the number, ordering and the thicknesses of the thin films. Differences between the perfectly designed mode of operation and its practical implementation may lead to side channels available to the eavesdropper. In our previous study [37], the WDM-components are found to have the transmittance windows outside the telecom range in agreement with the theoretical calculations reported in [47].
@@ III.2 Wavelength-division multiplexing components
Component
$\alpha_{\mathrm{min}}$ (dB)
$\lambda_{\mathrm{min}}$ (nm)
CWDM1
10.0
680
CWDM2
10.7
679
CWDM3
9.7
662
DWDM
10.3
665
@@ III.3 Isolating components
Owing to limitations of our experimental setup, we have not detected any noticeable transmission of light for backward propagation between adjacent (neighbouring) ports (for instance, propagation from the second port to the first port). In this case, the results are well below the noise level and bear close resemblance to the curves measured for the isolator.
IV Discussion and conclusion
In this paper, we have studied spectra of insertion losses of various fiber optical elements such as optical attenuators (see Fig. 2), WDM and isolating components (see Figs. 3 and 4, respectively) in the visible light range with wavelengths from $400$ nm to $800$ nm. Both the optical elements and the spectral range are important for understanding vulnerabilities of real-world QKD systems to hacking attacks whose efficiency being wavelength dependent may noticeably improve in the visible range. Our study is motivated by one of such attacks known as the induced-photorefraction attack [27, 45, 46].
The optical scheme behind our experimental technique enables us to perform measurements of the insertion losses with approximately $50$ dB dynamic range. It is simple to use and low-cost when compared with analogous schemes discussed, e.g., in Ref. [30]. An important point is that, by using suitably modified methods presented in Ref. [38], the proposed scheme can be improved so as to extend both the wavelength range and the dynamic range of measurements. As far as certification of QKD systems is concerned, this is a feasible and easy-to-use scheme that can be employed for characterization of optical elements insertion losses.
@@ IV Discussion and conclusion
Insertion losses for the light passing from the third port to the first port plotted against wavelength in Fig. 4(b) show that the well-known broadband filtering scheme [48] based on a Bragg grating connected to the second port of the circulator becomes ineffective in the 400-506 nm range. Thus it cannot be employed as a reliable tool to counter the IPA.
We summarize with the remark that the scattering based FOA characterized by nearly wavelength independent losses and utilized in combination with a couple of the isolating components will deliver at least 100 dB insertion losses which is enough to counter successful realization of the IPA.
@@ Acknowledgements.
The study is partially funded by the Ministry of Education and Science of the Russian Federation (Passport No. 2019-0903).

### v3

@@ IV.2 Wavelength-division multiplexing components
In QKD systems, wavelength-division multiplexing (WDM) components can be used as countermeasures by suppressing radiation with wavelengths outside the range utilized by the communication channel between Alice and Bob. In the simplest case, the key elements of the internal design of a fiber WDM component are thin films deposited on a substrate. Ideally, at the working wavelength, light will enter the common port (the com-port) and, after passing through the thin films and the substrate, will propagate to the corresponding port of the operating wavelengths grid. Other wavelengths are directed to the reflecting port (the ref-port). This port might be connected with a watchdog detector that monitor Eve’s probing pulses. The latter assumes that the sensitivity range of the watchdog detector is broad enough to cover wavelengths of potential attacks.
In practice, there are a number of factors limiting the efficiency of the WDM-components. These include the substrate transmittance as well as the number, ordering, and the thicknesses of the thin films. Differences between the perfectly designed mode of operation and its practical implementation may lead to side channels available to the eavesdropper. In our previous study [50], the WDM-components are found to have the transmittance windows outside the telecom range, in agreement with the theoretical calculations reported in [76].
@@
Component $\alpha_{\mathrm{min}}$ (dB) $\lambda_{\mathrm{min}}$ (nm)
CWDM1 10.0 680
CWDM2 10.7 679
CWDM3 9.7 662
DWDM 10.3 665
@@ IV.3 Isolating components
Owing to limitations of our experimental setup, we have not detected any noticeable transmission of light for backward propagation between adjacent (neighboring) ports (for instance, propagation from the second port to the first port). In this case, the results are well below the noise level and bear close resemblance to the curves measured for the isolator.
V IPA estimation and countermeasures
To show how the insertion losses influence security of QKD systems, we consider the four schemes described in Refs. [25, 35, 77, 33] (see Fig. 5 in Ref. [25], Fig. 1 in Ref. [35], Fig. 1 in Ref. [77], and Figs 2 and 3 in [33]). These schemes are ranged from the working prototype [77] to the ready-to-certification system [33, 35]. They serve as examples, QKD systems that differ in the stage of development. Since comprehensive information about the models and manufacturers of all the elements used in these schemes is not accessible, we have analyzed a variety of combinations involving different elements measured in previous sections (the results of additional measurements are presented in Appendix A.) and have chosen the data corresponding to the minimal and maximal total insertion losses. These data are shown in Fig. 7.
Figure 7: Estimated power, Eq. (7), that reaches the nearest investigated modulator in a system: (a) is for the scheme from [77], (b) is for the scheme from [25], (c) is for the SCW QKD scheme from [33], (d) is for the scheme from [35]; where blue (red) solid lines represents minimal (maximal) power available to an eavesdropper, black dashed line represents conservative mean of power needed for IPA realization: yellow, red and green triangles corresponding to estimations from [60], [59] and [28] respectively
In Fig. 7, the colored triangles indicate the powers and the corresponding wavelengths reported in Refs. [60, 59, 28] for experimentally confirmed successful IPA. Minimal value of these powers equals 3 nW [60]. This value can be regarded as a lower bound which is specified by the bold dashed line in Fig. 7. We consider the cases where the power exceeds this bound as vulnerabilities. Since the IPA efficiency is known to increase when the wavelength decreases, in our analysis, vulnerabilities in the short wavelength region are treated as the most important ones.
Figure 7a shows our data for the first scheme. It can be seen that, for the blue line, the power available to an eavesdropper is on average higher as compared to that for the red line, except for the two regions, where the red line is above the blue one. It is important that one of these regions is the range 400-410 nm, so that the red line indicates the power enough to realize successful IPA. We consider this range as the most dangerous case. From our data for the second scheme presented in Fig.7b we arrive at the conclusion that it is protected against IPA in all cases. Fig. 7d shows power estimations for different optical elements in SCW QKD scheme [33]. According to curves shown in Fig.7, the SCW QKD scheme is the most vulnerable to IPA system. Note that the noise level is determined by the insertion losses of the individual elements measured, with accuracy strongly affected when the value exceeds the dynamic measurement range. Individual outliers at certain wavelengths correspond to noise from individual pixels of the spectrometer. Therefore, the estimates can be considered conservative.
Countermeasure to this attack can be as follows. Before a connection of sender’s setup to an optical fiber link, it’s voltage, applied to the modulator, should be locked by either hardware or software (or both) and receiver’s setup should calibrate only. With the proposed approach to calibration, there is no chance that the power of prepared states will be higher than it should be – only lower. So again, if countermeasure is applied, IPA provides no advantage to an eavesdropper.
An alternative, straightforward countermeasure that can be implemented. The phenomenon of IPA requires a threshold optical power to emerge. Therefore, additional losses may be simply introduced in order to prevent illumination of the modulator with a power higher than a threshold.
VI Discussion and conclusion
In this work, we investigated security vulnerabilities introduced by induced photorefraction attacks across a range of QKD protocols and proposed countermeasures where applicable. Two principal categories of concern were identified. First, we examined protocols based on linearly independent quantum states, such as phase-encoded coherent-state and SCW QKD schemes. These protocols, unlike those based on linearly dependent states, are naturally resistant to IPA-enabled phase remapping. As demonstrated in Fig. 1, state overlap increases under phase distortion, which suppresses both unambiguous state discrimination and the success rate of any general measurement approach by an eavesdropper. The second class includes protocols that rely on amplitude or intensity modulators realized using phase modulators, which are commonly employed to control the signal power in decoy-state and SCW QKD systems. These systems remain vulnerable to IPA with partial energy recovery, emphasizing the need for careful system-level design and enhanced characterization of optical components, particularly in the visible spectral region. Thus, we have studied spectra of insertion losses of various fiber optical elements such as optical attenuators (see Fig. 4), WDM and isolating components (see Figs. 5 and 6, respectively) in the visible light range with wavelengths from $400$ nm to $800$ nm. Both the optical elements and the spectral range are important for understanding vulnerabilities of real-world QKD systems to hacking attacks, whose efficiency being wavelength dependent may noticeably improve in the visible range.
@@ VI Discussion and conclusion
Insertion losses for the light passing from the third port to the first port plotted against wavelength in Fig. 6(b) show that the well-known broadband filtering scheme [78] based on a Bragg grating connected to the second port of the circulator becomes ineffective in the 400-506 nm range. Thus, it cannot be employed as a reliable tool to counter the IPA.
Countermeasures against the IPA threat fall into two primary categories. The first is hardware-based and relatively straightforward: optical paths can be engineered to introduce additional insertion losses, thereby suppressing the intensity reaching the nearest lithium-niobate-based modulator. Figure 7 illustrates representative estimates of the incident intensity under various configurations. In practice, combining scattering-based FOAs with a sequence of isolators can yield more than 100 dB of loss—sufficient to block probe levels required for a successful IPA. The second category is algorithmic, focusing on voltage-locking mechanisms at the transmitter’s modulators. This technique prevents unauthorized modulation adjustments, ensuring that only a trusted receiver can calibrate system parameters and thereby reducing the overall attack surface.
@@ Acknowledgements.
B.N., A.I., A.G., A.T. and A.K. acknowledge the financial support of the Ministry of Science and Higher Education of the Russian Federation (No. FSER-2025–0007).
Appendix A Beamsplitters
Figure 8: Insertion losses measured in the 400-800 nm range for: (a) 50/50 three-port beamsplitter, (b) 50/50 four-ports beamsplitter, (c) three-port polarization beamsplitter, (d) 99/1 four-port beamsplitter, (e) and (f) illustrate ports numeration of beamsplitters
In order to present the calculated insertion losses for the investigated schemes in Fig. 7, we additionally have measured several beamsplitters. These beamsplitters are specified for $1550$ nm wavelength: $50/50$ beamsplitter with three connection ports, Fig. 8(a), $50/50$ beamsplitter with four connection ports, Fig. 8(b), one polarization beamsplitter with three ports, Fig. 8(c), and one $99/1$ with four ports, Fig. 8(d).
To simplify the perception of the figures, only the backward propagations as the insertion losses are presented, since they have a minor difference for both the forward and the backward propagation. Here we define forward connection as the propagation of scanning light from a port with a lower number to a port with a higher number, and backward connection as its opposite, for instance, see the box in Fig. 8.
For each beamsplitter, splitting ratios differ from the one of working wavelength, with an exception for the $99/1$ beamsplitter. This may be useful for an eavesdropper in the case when legitimate users have not measured insertion losses for beamsplitters within the considered range and have eventually chosen lower values. However, on the other hand, it may also be used as an additional countermeasure to attenuate injected light.